Write the slope-intercept equation for the line that passes through (-5, -8) and is perpendicular to 10x – 6y = -11 Please show all of your work. .

Answer by nerdybill(7384)

Write the slope-intercept equation for the line that passes through (-5, -8) and is perpendicular to 10x – 6y = -11 Please show all of your work. .
.
Determine slope of:
10x – 6y = -11
By rearranging into "slope-intercept" form:
10x – 6y = -11
10x = 6y -11
10x + 11 = 6y
(10/6)x + 11/6 = y
(5/3)x + 11/6 = y
slope is 5/3
a line perpendicular must have a "negative reciprocal":
(5/3)m = -1
m = -3/5 (slope of new line)
.
using the slope (-3/5) and (-5, -8) plug into the "point-slope" form:
y - y1 = m(x - x1)
y - (-8) = (-3/5)(x - (-5))
y + 8 = (-3/5)(x + 5)
y + 8 = (-3/5)x - 3
y = (-3/5)x - 11 (this is what they're looking for)